OpenStudy (darkigloo):

If dy/dt=-2y and if y=-1 when t=0, what is the value of t for which y=-1/2?

OpenStudy (science0229):

Take the antiderivative of both sides.

OpenStudy (darkigloo):

int (dy/y) = int(-2dt) so would that be lny=-2t+c?

OpenStudy (science0229):

When you take the antiderivative of both sides, it should look like\[\int\limits_{}^{}\frac{ dy }{ dt }=\int\limits_{}^{}-2ydt\]

OpenStudy (loser66):

I don't think it's the way to solve this problem

OpenStudy (science0229):

Correction: the left side should be \[\int\limits_{}^{}\frac{ dy }{ dt }dt\]

OpenStudy (science0229):

If you simplify that, it becomes y, and the right side simplifies to -2yt+C. So, the equation is y=-2yt+C

OpenStudy (loser66):

dy/dt = -2y it means y' + 2y =0 t =0 y = -1 mean y(0) =-1

OpenStudy (loser66):

solve for this you have y (t) = \(e^{-2t} +C\) replace y(0) = -1 you have C = -2 therefore, y (t) = \(e^{-2t} -2\) now calculate y(-1/2) = \(e^ {-1}-2 = \dfrac{1}{e} -2\)

OpenStudy (darkigloo):

how did you get y(t)=e^(-2t)+C

OpenStudy (loser66):

I guesss... hihihi... (joke) I believe that @amoodarya will give you the perfect explanation for that

OpenStudy (amoodarya):

\[\frac{ dy }{ dt }=-2y \\ \frac{ dy }{ y }=-2dt\\ \int\limits \frac{ dy }{ y }=\int\limits -2tdt\\ln y=-2t +c\\y=e^{-2t+c} \\y=e^{-2t} e^c=e^{-2t} a\]

OpenStudy (darkigloo):

where did the t come from in the third part?

OpenStudy (amoodarya):

i type with mistake there is not t in 3rd part

OpenStudy (darkigloo):

oh. so the integral of -2dt is -2t?

OpenStudy (amoodarya):

yes

OpenStudy (amoodarya):

like \[\int\limits -2dx=-2x+const\]

OpenStudy (darkigloo):

oh ok. and e^c is always a constant?

OpenStudy (amoodarya):

yes

OpenStudy (darkigloo):

so now how would I find the value of t for which y=-1/2?

OpenStudy (amoodarya):

like e^2 or e^3 ... or e^(ln5) =5

OpenStudy (darkigloo):

ok. i understand. how would I find the value of t for which y=-1/2? the answer choices are a) -ln2/2 b) -1/4 c) ln2/2 d)sqrt(2)/2 e) ln2

OpenStudy (amoodarya):

put t=0 and y=-1 y=a(e^(0)) -1=a *1 a=-1 so y(t)=-e^(-2t) y=-1/2 when -1/2=-1 e^(-2t) so e^(-2t)=1/2 ln e^(-2t)=ln 1/2 -2t=- ln 2 t=1/2 ln 2